OpenStudy (anonymous):

I need to check my answer! The table shows the values of y for each value of x: x y 2 1 3 1 4 -3 5 -3

OpenStudy (anonymous):

Part A: Does the table represent a relation that is a function? Justify your answer by using the values shown in the table. (2, 1) (3, 1) (4, -3) (5, -3) This is a function because no x-values or input values are the same.

OpenStudy (anonymous):

Part B: The function f(h) shown below represents Kylie's total earning from working in a cafe for h hours of work: f(h) = 15h + 5 f(h) = 15h + 5 f(8) = 15h + 5 15 * 8 = 120 120 + 5 = 125 f(8) = 125 Kylie made $125 for the number of hours she worked.

OpenStudy (anonymous):

Part C: Write an ordered pair to represent the input and output of the function in Part B when Kylie works for r hours. (8, 125)

OpenStudy (haseeb96):

for part A it is the function because function should satisfy the conditions that is all the values in domain ( here domain is x) are used and make only one image in range so here also all the terms in domain have only one range therefore it is the function

OpenStudy (anonymous):

ok, but what for part c?

OpenStudy (haseeb96):

yes u are correct for part C that is (8,125) because here when she spent 8 hours ( here time is domain or input ) then she get $125 (it is output or range ) Good job boyy

OpenStudy (anonymous):

Im a girl, but thanks! lol

OpenStudy (haseeb96):

okay but u are good in math